Two positive numbers have a difference of 8. The larger number is 2 less than 3 times the smaller number. Find the numbers.
Answer by macston(5194)

L=larger number; S=smaller number
L=3S-2
L-S=8 Substitute for L
(3S-2)-S=8 Add 2 to each side.
2S=10 Divide each side by 2
S=5 ANSWER 1: The smaller number is 5.
L-S=8
L-5=8 Add 5 to each side.
L=13 ANSWER 2: The larger number is 13.
CHECK:
L-2=3S
13+2=3(5)
15=15